1. Marinel received 1,450,500.00 pesos as her inheritance from her parents. She deposited the said amount in a time a deposit with 1% simple interest rate per annum, how much money will be accumulated after 7years?


1
Expert's answer
2021-11-29T01:47:15-0500

Simple interest


"A=P(1+rt)"

Given

"P=1450500"

"r=0.01"

"t=7"

Substitute


"A=1450500(1+0.01(7))=1552035"

1,552,035.00 pesos will be accumulated after 7years.